# A Metamorphic Testing Perspective on Knowledge Distillation for Language Models of Code: Does the Student Deeply Mimic the Teacher?

### v1

@@ 3.1.3 Adversarial Attack Approaches
In this study, adversarial attacks are employed to evaluate how deeply the student model mimics the teacher’s predictive behavior and internal representations during knowledge distillation. Adversarial attacks are typically categorized as black-box or white-box, depending on the attacker’s level of access to the model. Following the recommendation of Du et al. du2023extensive, we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code yang2022natural, zhang2020generating. The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions. To this end, we employ three established black-box techniques: ALERT yang2022natural, Metropolis Hastings Modifier (MHM) zhang2020generating, and WIR-Random zeng2022extensive, each of which performs identifier substitution to assess model robustness without altering code functionality. ALERT yang2022natural considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier. MHM zhang2020generating utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration. WIR-Random zeng2022extensive ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates. It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications yang2022natural, zhang2020generating.
@@ 3.1.4 Experimental Settings
We fine-tune CodeBERT using the hyperparameter settings and training pipeline provided by the CodeXGLUE benchmark lu2021codexglue, while for GraphCodeBERT, we adopt the settings specified in its original paper guo2020graphcodebert. To ensure consistency across our experiments, we adopt the data-splitting strategy used in knowledge distillation–based approaches such as MORPH and AVATAR, which divide the training data into two equal halves: one for fine-tuning and the other for knowledge distillation. We apply the same fine-tuning split to all models. Despite being trained on only half of the original dataset, all models achieve performance comparable to the state-of-the-art results reported in the CodeXGLUE benchmark lu2021codexglue, the GraphCodeBERT paper guo2020graphcodebert, and knowledge distillation–based compression studies such as Compressor, AVATAR, and MORPH.
@@ 3.1.5 Evaluation Metrics
Following the evaluation protocols of CodeXGLUE lu2021codexglue, Yang et al. yang2022natural, and Shi et al. shi2022compressing, shi2024greening, we use accuracy as the primary metric to evaluate model performance on downstream tasks.
To evaluate the performance of both teacher and student models under adversarial conditions, we measure the effectiveness of the applied attack methods. A higher attack effectiveness indicates greater model vulnerability to adversarial perturbations. We employ the Attack Success Rate (ASR) metric yang2022natural to measure the effectiveness of an attack method. ASR measures the proportion of adversarial samples for which the model fails to produce a correct prediction after the attack and is formally defined as follows:
@@
The robustness of compressed models to adversarial attacks depends strongly on the quality of the adversarial examples generated. If these examples contain syntactic or semantic errors, model mispredictions may not accurately reflect a lack of robustness but rather the model’s inability to handle unrealistic inputs. Moreover, excessive modifications to the original code during adversarial example generation can lead to out-of-distribution instances, thereby reducing the reliability of robustness assessments. To ensure a fair and meaningful evaluation, we conduct a qualitative analysis to assess the effectiveness of the generated adversarial examples and to verify that the applied perturbations are minimal yet sufficient to mislead the models. Because the inputs for the selected software analytics tasks are source code, we employ four established metrics from Du et al. du2023extensive: Identifier Change Rate (ICR), Token Change Rate (TCR), Average Edit Distance (AED), and Average Code Similarity (ACS), to measure the quality of the adversarial examples.
@@ 3.2 Empirical Results
Overall, the results indicate that these attack techniques effectively preserve the syntactic integrity of the generated adversarial examples. Specifically, the examples consistently exhibit high ACS values and low TCR scores across both CodeBERT and GraphCodeBERT models, demonstrating minimal syntactical distortion. The adversarial examples also maintain strong overall quality across the AED and ICR metrics for both the teacher and student models. Although a slight degradation is observed in certain cases—for instance, the MHM attack shows comparatively higher AED values—the overall differences remain limited. We further perform the Friedman test friedman1937use to determine whether the three attack techniques (ALERT, MHM, and WIR-Random) generate adversarial examples of significantly different quality for each model variant of CodeBERT and GraphCodeBERT. A $p$-value below 0.05 indicates that the differences in the quality of adversarial examples produced by the three attack techniques are statistically significant. However, the experimental results show no statistically significant differences among the attack methods for either CodeBERT ($p=0.47$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively) or GraphCodeBERT ($p=0.63$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively). These results indicate that all three attack techniques generate adversarial examples of comparable quality across the evaluated metrics (ICR, TCR, AED, and ACS).
Although the preceding results show that the adversarial examples essentially preserve syntactic and semantic quality across both teacher and student models, it remains important to examine whether any subtle quality differences exist between these two model categories. To investigate this, we compare the quality of adversarial examples generated by each attack technique for the teacher and student models across all the evaluation metrics. The Wilcoxon signed-rank test woolson2007wilcoxon is used to assess the statistical significance of these differences. The comparison between the Teacher and AVATAR models ($p=0.13$) and between the Teacher and MORPH models ($p=0.85$) indicates that the quality of adversarial examples remains statistically comparable, with no significant differences observed in either case. These findings suggest that the attack techniques generate adversarial examples of similar syntactic and semantic integrity across both teacher and student models, confirming that the subsequent robustness analysis is based on adversarial inputs of consistent quality and that any observed performance degradation can be attributed to the models’ inability to maintain robustness against adversarial perturbations.
@@ 4 Approach: MetaCompress Framework
The empirical findings in Section 3.2 show that traditional accuracy-based evaluation is insufficient to capture behavioral discrepancies between teacher and student models, which become evident under adversarial conditions due to the student models’ inability to deeply mimic their teacher counterparts’ predictive behavior and internal representations. Motivated by this observation, we introduce MetaCompress, a metamorphic testing framework for evaluating the behavioral fidelity of compressed models by framing the task as a classical software testing problem and applying a metamorphic testing strategy. The framework identifies key behavioral invariants that should be preserved after model compression and formalizes a set of output-based metamorphic relations that serve as oracle-free tests to assess behavioral fidelity between teacher and student models. Figure 3 presents the overall design and workflow of the proposed MetaCompress framework.
@@ 4.1 Behavioral Fidelity as a Software Testing Problem
In traditional software engineering, techniques such as regression testing wong1997study, mutation testing jia2010analysis, and metamorphic testing chen2018metamorphic are commonly employed to assess the reliability of modified systems. However, evaluating language models of code presents unique challenges compared to traditional software systems. For example, traditional software typically produces deterministic outputs, while language models of code yield probabilistic outputs, and knowledge distillation may alter output distributions even when top-1 predictions remain unchanged. Additionally, we aim to assess the behavioral fidelity of a student model compared to its teacher counterpart. Although ground-truth labels are available for tasks such as clone detection or vulnerability prediction, our objective is not to directly compare the outputs of the two models against the ground-truth. Instead, we compare their outputs after model execution, treating the teacher model as the behavioral reference. Notably, the teacher model’s behavior is not known until it is executed on test data. We refer to this scenario as an instance of the oracle problem, where the absence of a definitive oracle hinders straightforward correctness verification.
As a result, conventional software testing techniques such as regression testing, which depend on exact output matches, are often too rigid—minor differences in probabilistic outputs or generated sequences do not necessarily indicate functional errors. Mutation testing is also ill-suited, as neither model is inherently faulty. In contrast, metamorphic testing provides a principled approach to overcoming the oracle problem by defining expected relations (as detailed in Section 4.2) between the outputs of the teacher and student models, without requiring explicit ground-truth. These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs. Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models xie2020mettle, sun2024fairness, kwan2024mt, zhang2020machine, wu2025detecting, yang2025hallucination, demonstrating its versatility and effectiveness in oracle-free evaluation scenarios. Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
@@ 4.2.2 MR2 - Probability Distribution Similarity
To quantify the difference between the two probability distributions, we use the Kullback–Leibler divergence, as defined in “(1)”. Given a tolerance $\delta=0.5$ (Inspired by sigmoid-based methods for binary classification lecun2015deep), we then count the test instances for which MR2 holds:
@@ 4.2.4 MR4 - Calibration Alignment
Top-1 agreement alone does not indicate whether the student model is well-calibrated. Poor calibration, where predicted probabilities do not reflect true likelihoods, can lead to over- or under-confidence, which adversely affects downstream decision-making. Prior studies shang2024enhancing, ji2024beware on vision and language models demonstrate that compression techniques such as quantization can significantly distort model calibration, even when accuracy remains unaffected. These findings motivate the design of this MR to assess behavioral fidelity through the lens of calibration alignment.
@@ 5 Discussion
Ba et al. ba2014deep stated that ”If the student model learns to mimic the teacher model perfectly, it makes exactly the same predictions and mistakes as the teacher model.” However, achieving such mimicry is extremely challenging when there exists a substantial gap in network capacity wang2021knowledge. For instance, the knowledge distillation technique Compressor reduces the size of the CodeBERT model from 476 MB to only 3 MB. Consequently, although the student models produced by Compressor, AVATAR, and MORPH achieve accuracy comparable to the teacher, their performance deteriorates considerably under adversarial attacks. Even though MORPH attempts to distill knowledge more robustly by incorporating metamorphic attacks through metamorphic code transformations, the resulting student models still exhibit notable adversarial vulnerabilities. These observations indicate that the knowledge distillation strategies employed by Compressor, AVATAR, and MORPH do not enable the student models to deeply mimic the teacher’s behavior, leaving a substantial knowledge gap in behavioral fidelity between the teacher and the student models.
@@
    Wang et al. wang2025empirical empirically demonstrated that the same student model distilled from a teacher performs significantly better than when trained using the standard fine-tuning approach on the same task. Since the teacher model typically possesses much greater network capacity, it can learn more complex decision boundaries and have greater robustness against adversarial attacks. Minimizing the knowledge gap, therefore, enables the distilled student model to better approximate these boundaries during distillation, potentially enhancing its resilience to adversarial attacks.
@@ 6.1 Internal Validity
A potential threat to internal validity may arise if differences in performance between the student and teacher models are influenced by inconsistent training procedures, fine-tuning configurations, or data handling, rather than the knowledge distillation approaches themselves. To mitigate this threat, we applied uniform training protocols and hyperparameter settings across all models, both student and teacher, following the configurations reported in prior knowledge distillation studies shi2022compressing, shi2024greening, panichella2025metamorphic. Additionally, we used identical data splits for fine-tuning and evaluation to ensure consistency and fairness in comparison. Maintaining such a consistent experimental setup enhances the internal validity of our findings by ensuring that the observed effects are due to the student model’s inability to deeply mimic the teacher model’s behavior, rather than differences in hyperparameters or data partitions. Additionally, the experimental results may be influenced by potential errors in implementing the training pipelines for both student and teacher models. To mitigate this threat, we adopted the standardized training procedures provided by the CodeXGLUE benchmark for teacher model fine-tuning and evaluation. For knowledge distillation, we utilized publicly available, off-the-shelf student models obtained from Compressor [^8] https://github.com/soarsmu/Compressor, AVATAR [^9] https://github.com/soarsmu/Avatar, and MORPH [^10] https://github.com/apanichella/morph-distillation/tree/main.
@@ 7 Related Work
A substantial body of research on model compression has emerged in the fields of computer vision and natural language processing (NLP), aiming to develop large language models that are energy efficient, environmentally sustainable, and optimized for memory usage xu2023survey, zhu2024survey, xu2025resource, xu2021survey, gordon2020compressing, xu2021beyond, ye2019adversarial, du2021robustness, gourtani2024improving, zhu2022safety, sanh2019distilbert, sun2019patient, jiao2019tinybert, buciluǎ2006model, jiang2023lion, zhang2018structadmm, tang2019distilling, xu2020bert. To this end, several techniques, including pruning sanh2020movement, quantization zafrir2019q8bert, and knowledge distillation (KD) hinton2015distilling, have been widely explored. More recently, these compression techniques have been extended to software analytics to improve the efficiency of language models of code. Wei et al. wei2023towards evaluated quantized models on code generation tasks, examining accuracy, resource consumption, and environmental impact, and found that quantization can substantially improve model efficiency with minimal degradation in accuracy or robustness when applied under suitable conditions. Saad et al. saad2024alpine introduced ALPINE, a pruning method that is independent of programming language and effectively reduces the computational cost of code-language models. Aloisio et al. d2024compression conducted an empirical study on how different compression strategies, such as pruning, quantization, and knowledge distillation, affect the performance and resource efficiency of the CodeBERT model across three software analytics tasks: code summarization, vulnerability detection, and code search. Afrin et al. afrin2025quantization empirically examined the effects of quantizing large code models beyond functional correctness, focusing on qualitative aspects and static properties of the generated code. While pruning and quantization improve efficiency, their compression capacity is limited, as excessive compression often harms model performance. Hence, this study focuses on knowledge distillation, which enables extremely compact models without compromising significant predictive performance.
@@
Shi et al. shi2022compressing proposed Compressor, a KD-based method evaluated on CodeBERT and GraphCodeBERT for two code understanding tasks: vulnerability prediction and clone detection. Experimental results show that Compressor significantly reduces inference time and model size, down to 3 MB, while preserving model effectiveness. In a follow-up study, Shi et al. shi2024greening introduced AVATAR, another KD-based approach designed to mitigate energy consumption and carbon emissions, while also maintaining model effectiveness and reducing inference latency. Building on AVATAR, MORPH panichella2025metamorphic was proposed as a KD method that combines metamorphic testing with many-objective optimization to enhance model robustness on metamorphic code, as well as accuracy, size, and efficiency. Experiments on CodeBERT and GraphCodeBERT showed that MORPH achieved up to 73% greater robustness and 38% greater efficiency than AVATAR, without sacrificing accuracy or increasing model size. However, in MORPH, the metamorphic attack using a metamorphic sample technically differs from the adversarial attack using an adversarial example. Metamorphic attacks are performed only once on the metamorphic sample, whereas adversarial attacks are performed by repeatedly querying the model, modifying the original sample to assess the model’s actual robustness. Our empirical evaluation also reveals that the student models obtained from MORPH experience greater performance degradation under adversarial attacks than under metamorphic attacks, compared to the student models obtained from AVATAR.
@@
Recently, Liu et al. liu2026pioneer proposed a KD-based approach, PIONEER, that enhances the robustness of compressed code models while preserving accuracy. Using a solver-constrained genetic algorithm and robustness evaluation with ALERT-generated examples, it achieved an 87.54% robustness gain with only a 1.67% drop in accuracy across CodeBERT, GraphCodeBERT, and CodeT5. Recently, Wang et al. wang2025empirical empirically evaluated the effectiveness of logit-based and feature-based KD approaches for code understanding tasks. Their study showed that KD consistently outperforms fine-tuning on code understanding tasks, with feature-based methods and code-specific teachers yielding the best results.

### v2

@@ 3.1.3 Adversarial Attack Approaches
In this study, adversarial attacks are employed to evaluate how deeply the student model mimics the teacher’s predictive behavior and internal representations during knowledge distillation. Adversarial attacks are typically categorized as black-box or white-box, depending on the attacker’s level of access to the model. Following the recommendation of Du et al. Du et al. [2023], we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code Yang et al. [2022], Zhang et al. [2020a]. The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions. To this end, we employ three established black-box techniques: ALERT Yang et al. [2022], Metropolis Hastings Modifier (MHM) Zhang et al. [2020a], and WIR-Random Zeng et al. [2022], each of which performs identifier substitution to assess model robustness without altering code functionality. ALERT Yang et al. [2022] considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier. MHM Zhang et al. [2020a] utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration. WIR-Random Zeng et al. [2022] ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates. It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications Yang et al. [2022], Zhang et al. [2020a].
@@ 3.1.4 Experimental Settings
We fine-tune CodeBERT using the hyperparameter settings and training pipeline provided by the CodeXGLUE benchmark Lu et al. [2021], while for GraphCodeBERT, we adopt the settings specified in its original paper Guo et al. [2020]. To ensure consistency across our experiments, we adopt the data-splitting strategy used in knowledge distillation–based approaches such as MORPH and AVATAR, which divide the training data into two equal halves: one for fine-tuning and the other for knowledge distillation. We apply the same fine-tuning split to all models. Despite being trained on only half of the original dataset, all models achieve performance comparable to the state-of-the-art results reported in the CodeXGLUE benchmark Lu et al. [2021], the GraphCodeBERT paper Guo et al. [2020], and knowledge distillation–based compression studies such as Compressor, AVATAR, and MORPH.
@@ 3.1.5 Evaluation Metrics
Following the evaluation protocols of CodeXGLUE Lu et al. [2021], Yang et al. Yang et al. [2022], and Shi et al. Shi et al. [2023, 2024a], we use accuracy as the primary metric to evaluate model performance on downstream tasks.
To evaluate the performance of both teacher and student models under adversarial conditions, we measure the effectiveness of the applied attack methods. A higher attack effectiveness indicates greater model vulnerability to adversarial perturbations. We employ the Attack Success Rate (ASR) metric Yang et al. [2022] to measure the effectiveness of an attack method. ASR measures the proportion of adversarial samples for which the model fails to produce a correct prediction after the attack and is formally defined as follows:
@@
The robustness of compressed models to adversarial attacks depends strongly on the quality of the adversarial examples generated. If these examples contain syntactic or semantic errors, model mispredictions may not accurately reflect a lack of robustness but rather the model’s inability to handle unrealistic inputs. Moreover, excessive modifications to the original code during adversarial example generation can lead to out-of-distribution instances, thereby reducing the reliability of robustness assessments. To ensure a fair and meaningful evaluation, we conduct a qualitative analysis to assess the effectiveness of the generated adversarial examples and to verify that the applied perturbations are minimal yet sufficient to mislead the models. Because the inputs for the selected software analytics tasks are source code, we employ four established metrics from Du et al. Du et al. [2023]: Identifier Change Rate (ICR), Token Change Rate (TCR), Average Edit Distance (AED), and Average Code Similarity (ACS), to measure the quality of the adversarial examples.
@@ 3.2 Empirical Results
Overall, the results indicate that these attack techniques effectively preserve the syntactic integrity of the generated adversarial examples. Specifically, the examples consistently exhibit high ACS values and low TCR scores across both CodeBERT and GraphCodeBERT models, demonstrating minimal syntactical distortion. The adversarial examples also maintain strong overall quality across the AED and ICR metrics for both the teacher and student models. Although a slight degradation is observed in certain cases—for instance, the MHM attack shows comparatively higher AED values—the overall differences remain limited. We further perform the Friedman test Friedman [1937] to determine whether the three attack techniques (ALERT, MHM, and WIR-Random) generate adversarial examples of significantly different quality for each model variant of CodeBERT and GraphCodeBERT. A $p$-value below 0.05 indicates that the differences in the quality of adversarial examples produced by the three attack techniques are statistically significant. However, the experimental results show no statistically significant differences among the attack methods for either CodeBERT ($p=0.47$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively) or GraphCodeBERT ($p=0.63$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively). These results indicate that all three attack techniques generate adversarial examples of comparable quality across the evaluated metrics (ICR, TCR, AED, and ACS).
Although the preceding results show that the adversarial examples essentially preserve syntactic and semantic quality across both teacher and student models, it remains important to examine whether any subtle quality differences exist between these two model categories. To investigate this, we compare the quality of adversarial examples generated by each attack technique for the teacher and student models across all the evaluation metrics. The Wilcoxon signed-rank test Woolson [2007] is used to assess the statistical significance of these differences. The comparison between the Teacher and AVATAR models ($p=0.13$) and between the Teacher and MORPH models ($p=0.85$) indicates that the quality of adversarial examples remains statistically comparable, with no significant differences observed in either case. These findings suggest that the attack techniques generate adversarial examples of similar syntactic and semantic integrity across both teacher and student models, confirming that the subsequent robustness analysis is based on adversarial inputs of consistent quality and that any observed performance degradation can be attributed to the models’ inability to maintain robustness against adversarial perturbations.
@@ 4 Approach: MetaCompress Framework
The empirical findings in Section 3.2 show that traditional accuracy-based evaluation is insufficient to capture behavioral discrepancies between teacher and student models, which become evident under adversarial conditions due to the student models’ inability to deeply mimic their teacher counterparts’ predictive behavior and internal representations. Motivated by this observation, we introduce MetaCompress, a metamorphic testing framework for evaluating the behavioral fidelity of compressed models by framing the task as a classical software testing problem and applying a metamorphic testing strategy. The framework identifies key behavioral invariants that should be preserved after knowledge distillation and formalizes a set of output-based metamorphic relations that serve as oracle-free tests to assess behavioral fidelity between teacher and student models. Figure 3 presents the overall design and workflow of the proposed MetaCompress framework.
@@ 4.1 Behavioral Fidelity as a Software Testing Problem
In traditional software engineering, techniques such as regression testing Wong et al. [1997], mutation testing Jia and Harman [2010], and metamorphic testing Chen et al. [2018] are commonly employed to assess the reliability of modified systems. However, evaluating language models of code presents unique challenges compared to traditional software systems. For example, traditional software typically produces deterministic outputs, while language models of code yield probabilistic outputs, and knowledge distillation may alter output distributions even when top-1 predictions remain unchanged. Additionally, we aim to assess the behavioral fidelity of a student model compared to its teacher counterpart. Although ground-truth labels are available for tasks such as clone detection or vulnerability prediction, our objective is not to directly compare the outputs of the two models against the ground-truth. Instead, we compare their outputs after model execution, treating the teacher model as the behavioral reference. Notably, the teacher model’s behavior is not known until it is executed on test data. We refer to this scenario as an instance of the oracle problem, where the absence of a definitive oracle hinders straightforward correctness verification.
As a result, conventional software testing techniques such as regression testing, which depend on exact output matches, are often too rigid—minor differences in probabilistic outputs or generated sequences do not necessarily indicate functional errors. Mutation testing is also ill-suited, as neither model is inherently faulty. In contrast, metamorphic testing provides a principled approach to overcoming the oracle problem by defining expected relations (as detailed in Section 4.2) between the outputs of the teacher and student models, without requiring explicit ground-truth. These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs. Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models Xie et al. [2020], Sun et al. [2024], Kwan et al. [2024], Zhang et al. [2020b], Wu et al. [2025], Yang et al. [2025], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios. Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
@@ 4.2.2 MR2 - Probability Distribution Similarity
To quantify the difference between the two probability distributions, we use the Kullback–Leibler divergence, as defined in “(1)”. Given a tolerance $\delta=0.5$ (Inspired by sigmoid-based methods for binary classification LeCun et al. [2015]), we then count the test instances for which MR2 holds:
@@ 4.2.4 MR4 - Calibration Alignment
Top-1 agreement alone does not indicate whether the student model is well-calibrated. Poor calibration, where predicted probabilities do not reflect true likelihoods, can lead to over- or under-confidence, which adversely affects downstream decision-making. Prior studies Shang et al. [2024], Ji et al. [2024] on vision and language models demonstrate that compression techniques such as quantization can significantly distort model calibration, even when accuracy remains unaffected. These findings motivate the design of this MR to assess behavioral fidelity through the lens of calibration alignment.
@@ 5 Discussion
Ba et al. Ba and Caruana [2014] stated that ”If the student model learns to mimic the teacher model perfectly, it makes exactly the same predictions and mistakes as the teacher model.” However, achieving such mimicry is extremely challenging when there exists a substantial gap in network capacity Wang and Yoon [2021]. For instance, the knowledge distillation technique Compressor reduces the size of the CodeBERT model from 476 MB to only 3 MB. Consequently, although the student models produced by Compressor, AVATAR, and MORPH achieve accuracy comparable to the teacher, their performance deteriorates considerably under adversarial attacks. Even though MORPH attempts to distill knowledge more robustly by incorporating metamorphic attacks through metamorphic code transformations, the resulting student models still exhibit notable adversarial vulnerabilities. These observations indicate that the knowledge distillation strategies employed by Compressor, AVATAR, and MORPH do not enable the student models to deeply mimic the teacher’s behavior, leaving a substantial knowledge gap in behavioral fidelity between the teacher and the student models.
@@
    Wang et al. Wang et al. [2025] empirically demonstrated that the same student model distilled from a teacher performs significantly better than when trained using the standard fine-tuning approach on the same task. Since the teacher model typically possesses much greater network capacity, it can learn more complex decision boundaries and have greater robustness against adversarial attacks. Minimizing the knowledge gap, therefore, enables the distilled student model to better approximate these boundaries during distillation, potentially enhancing its resilience to adversarial attacks.
@@ 6.1 Internal Validity
A potential threat to internal validity may arise if differences in performance between the student and teacher models are influenced by inconsistent training procedures, fine-tuning configurations, or data handling, rather than the knowledge distillation approaches themselves. To mitigate this threat, we applied uniform training protocols and hyperparameter settings across all models, both student and teacher, following the configurations reported in prior knowledge distillation studies Shi et al. [2023, 2024a], Panichella [2025]. Additionally, we used identical data splits for fine-tuning and evaluation to ensure consistency and fairness in comparison. Maintaining such a consistent experimental setup enhances the internal validity of our findings by ensuring that the observed effects are due to the student model’s inability to deeply mimic the teacher model’s behavior, rather than differences in hyperparameters or data partitions. Additionally, the experimental results may be influenced by potential errors in implementing the training pipelines for both student and teacher models. To mitigate this threat, we adopted the standardized training procedures provided by the CodeXGLUE benchmark for teacher model fine-tuning and evaluation. For knowledge distillation, we utilized publicly available, off-the-shelf student models obtained from Compressor [^8] https://github.com/soarsmu/Compressor, AVATAR [^9] https://github.com/soarsmu/Avatar, and MORPH [^10] https://github.com/apanichella/morph-distillation/tree/main.
@@ 7 Related Work
A substantial body of research on model compression has emerged in the fields of computer vision and natural language processing (NLP), aiming to develop large language models that are energy efficient, environmentally sustainable, and optimized for memory usage Xu and McAuley [2023], Zhu et al. [2024], Xu et al. [2025, 2021b], Gordon et al. [2020], Xu et al. [2021a], Ye et al. [2019], Du et al. [2021], Gourtani and Meratnia [2024], Zhu et al. [2022], Sanh [2019], Sun et al. [2019], Jiao et al. [2019], buciluǎ2006model, Jiang et al. [2023], Zhang et al. [2018], Tang et al. [2019], Xu et al. [2020]. To this end, several techniques, including pruning Sanh et al. [2020], quantization Zafrir et al. [2019], and knowledge distillation (KD) Hinton et al. [2015], have been widely explored. More recently, these compression techniques have been extended to software analytics to improve the efficiency of language models of code. Wei et al. Wei et al. [2023] evaluated quantized models on code generation tasks, examining accuracy, resource consumption, and environmental impact, and found that quantization can substantially improve model efficiency with minimal degradation in accuracy or robustness when applied under suitable conditions. Saad et al. Saad et al. [2024] introduced ALPINE, a pruning method that is independent of programming language and effectively reduces the computational cost of code-language models. Aloisio et al. d’Aloisio et al. [2024] conducted an empirical study on how different compression strategies, such as pruning, quantization, and knowledge distillation, affect the performance and resource efficiency of the CodeBERT model across three software analytics tasks: code summarization, vulnerability detection, and code search. Afrin et al. Afrin et al. [2025] empirically examined the effects of quantizing large code models beyond functional correctness, focusing on qualitative aspects and static properties of the generated code. While pruning and quantization improve efficiency, their compression capacity is limited, as excessive compression often harms model performance. Hence, this study focuses on knowledge distillation, which enables extremely compact models without compromising significant predictive performance.
@@
Shi et al. Shi et al. [2023] proposed Compressor, a KD-based method evaluated on CodeBERT and GraphCodeBERT for two code understanding tasks: vulnerability prediction and clone detection. Experimental results show that Compressor significantly reduces inference time and model size, down to 3 MB, while preserving model effectiveness. In a follow-up study, Shi et al. Shi et al. [2024a] introduced AVATAR, another KD-based approach designed to mitigate energy consumption and carbon emissions, while also maintaining model effectiveness and reducing inference latency. Building on AVATAR, MORPH Panichella [2025] was proposed as a KD method that combines metamorphic testing with many-objective optimization to enhance model robustness on metamorphic code, as well as accuracy, size, and efficiency. Experiments on CodeBERT and GraphCodeBERT showed that MORPH achieved up to 73% greater robustness and 38% greater efficiency than AVATAR, without sacrificing accuracy or increasing model size. However, in MORPH, the metamorphic attack using a metamorphic sample technically differs from the adversarial attack using an adversarial example. Metamorphic attacks are performed only once on the metamorphic sample, whereas adversarial attacks are performed by repeatedly querying the model, modifying the original sample to assess the model’s actual robustness. Our empirical evaluation also reveals that the student models obtained from MORPH experience greater performance degradation under adversarial attacks than under metamorphic attacks, compared to the student models obtained from AVATAR.
@@
Recently, Liu et al. Liu et al. [2026] proposed a KD-based approach, PIONEER, that enhances the robustness of compressed code models while preserving accuracy. Using a solver-constrained genetic algorithm and robustness evaluation with ALERT-generated examples, it achieved an 87.54% robustness gain with only a 1.67% drop in accuracy across CodeBERT, GraphCodeBERT, and CodeT5. Recently, Wang et al. Wang et al. [2025] empirically evaluated the effectiveness of logit-based and feature-based KD approaches for code understanding tasks. Their study showed that KD consistently outperforms fine-tuning on code understanding tasks, with feature-based methods and code-specific teachers yielding the best results.